OpenStudy (anonymous):

Converge or Diverge (With justification). Series from 1 to infinity of ((1+n)^.5-(n-1)^.5)/n

OpenStudy (zarkon):

multiply top and bottom by \[\sqrt{1+n}+\sqrt{n-1}\]

OpenStudy (zarkon):

then simplify...that should lead you in the right direction

OpenStudy (anonymous):

2/n(1+n)^.5-(n-1)^.5)... so does that converge?

OpenStudy (zarkon):

you get \[\sum_{n=1}^{\infty}\frac{2}{n\left(\sqrt{1+n}+\sqrt{n-1}\right)}\]

OpenStudy (anonymous):

I realize, but how do i justify if it converges or diverges?

OpenStudy (zarkon):

use the limit comparison test

OpenStudy (anonymous):

Thank you.